# Faster Quantum Simulation Of Markovian Open Quantum Systems Via Randomisation

### v2

@@ Appendix A Proofs Of Theorems and Lemmas In Section 2
$=\leq\left\|T\right\|_{\diamond}\left\|T-V\right\|_{\diamond}+\left\|T-V\right\|_{\diamond}\left\|V\right\|_{\diamond}$ (A.1)
@@
(of Theorem 1.) To begin the proof, we define the parameter $\tau=t/N$ to be a small time step, and we recognise that
@@
where the last inequality is obtained by the fact that ^$\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}\leq\Lambda$ for all $k=1,...,M$. To complete the bound we must compute the restricted sum in equation (A.21). We make use of Lemma 1 found in Appendix B to do this which leads to,
@@
(proof of Theorem 2.) Theorem 2 is proved in a similar manner to Theorem 1. Start by defining $\tau=t/N$ and applying Lemma 1. to equation (24),
@@
We now need to bound the diamond norm of the product of the summands ^$\hat{\mathcal{L}}_{k}$ in equation (A.37). This is done using submultiplicativity of the diamond norm as
@@
where the last inequality is obtained using the fact that ^$\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}\leq\Lambda$ for all $k$. Using the inquality (A.40) with equation (A.37) yields

### v3

@@ Appendix A Proofs Of Theorems and Lemmas In Section 2
$\leq\left\|T\right\|_{\diamond}\left\|T-V\right\|_{\diamond}+\left\|T-V\right\|_{\diamond}\left\|V\right\|_{\diamond}$ (A.1)
@@
(Theorem 1.) To begin the proof, we define the parameter $\tau=t/N$ to be a small time step, and we recognise that
@@
where the last inequality is obtained by the fact that $\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}\leq\Lambda$ for all k=1,,Mk=1,...,M. To complete the bound we must compute the restricted sum in equation (A.21). We make use of Lemma B.1 found in Appendix B to do this which leads to,
@@
(Theorem 2.) Theorem 2 is proved in a similar manner to Theorem 1. Start by defining $\tau=t/N$ and applying Lemma 1. to equation (24),
@@
We now need to bound the diamond norm of the product of the summands $\hat{\mathcal{L}}_{k}$ in equation (A.37). This is done using submultiplicativity of the diamond norm as
@@
where the last inequality is obtained using the fact that $\left\|\hat{\mathcal{L}}_{k}\right\|_{\diamond}\leq\Lambda$ for all $k$. Using the inquality (A.40) with equation (A.37) yields